# The Binary Fraction of Stars in the Dwarf Galaxy Ursa Minor via Dark Energy Spectroscopic Instrument

### v1

@@ Abstract
journal: ApJthanks: qtskyxt@sjtu.edu.cnthanks: wenting.wang@sjtu.edu.cnthanks: Sergey.Koposov@ed.ac.uk
1 Introduction
@@
Binary stars play a crucial role in our understanding of the star formation process, galactic dynamics, and the evolution history of galaxies (2001ApJ...556..265W; 2013ARA&A..51..269D; 2018ApJ...854..147B). Binaries are common in our Milky Way Galaxy, with studies showing that binary systems are identified more often than single stars in the solar neighborhood (1969JRASC..63..275H; 1991A&A...248..485D; 2010ApJS..190....1R). These earlier studies mostly focus on the statistical properties of binary stars within a limited range, typically extending only to about 30 pc from the Sun.
@@
With advances in observational techniques, investigations of binary populations have extended (2018AJ....156...18P) beyond the solar neighborhood to open clusters (2002AJ....123.1570P; 2020ApJ...901...49L), globular clusters (2007MNRAS.380..781S; 2012A&A...540A..16M) and dwarf galaxies (2010ApJ...722L.209M; 2013ApJ...779..116M). In dense stellar environments, binaries are likely to have undergone significant dynamical processes such as stellar encounters (2011A&A...528A.144K; 2020MNRAS.496.5176D), which can disrupt binary systems. Compared with globular clusters, dwarf galaxies are more diffuse, making close stellar encounters among their member stars less frequent. The lower stellar densities in dwarf galaxies may allow a higher fraction of binaries to survive than those in dense globular clusters.
@@
Most of the classical dwarf galaxies around our Milky Way are distant, and thus binaries in dwarf galaxies are difficult to resolve directly. The influence of binaries is often inferred from their contribution to observed line-of-sight velocity (LOSV) variations. For example, binary orbital motions can inflate the measured velocity dispersion of a galaxy, thereby complicating dynamical mass estimates (1996AJ....111..750O; 1996MNRAS.279..108H; 2010ApJ...722L.209M; 2011ApJ...736..146K; 2019MNRAS.487.2961M; 2023ApJ...956...91W). With advances in spectroscopic surveys, the number of stars observed in dwarf galaxies has increased substantially, from less than a hundred in earlier studies (1995A&A...300...31Q; 1995AJ....110.2131A; 1995AJ....109..151V; 2002MNRAS.330..792K; 2003ApJ...588L..21K) to several hundred or even over a thousand in those known dwarf galaxies such as Carina, Fornax, Sculptor, Sextans, etc (2009AJ....137.3100W; 2023ApJS..268...19W).
@@
The growing dataset has enabled several studies to develop statistical methods for estimating binary fractions by modeling the inflation of velocity dispersions caused by unresolved binaries. As described in 2010ApJ...721.1142M, a method was introduced to estimate the binary fraction in dwarf galaxies based on their velocity dispersion. 2011ApJ...733...46S applied this approach to Segue 1 to determine the velocity dispersion, removing the influence from the binaries. A slightly later study by 2013ApJ...779..116M conducted a detailed binary analysis of several dwarf galaxies, including Carina, Fornax, etc. Combining several previous datasets, 2017ApJ...836..202S compiled multi-epoch LOSV measurements and applied a newly developed method to study the binary population in Leo II. A more recent study by 2018AJ....156..257S (hereafter 2018AJ....156..257S) further extended this methodology by combining LOSV measurements from multiple sources collected over approximately 30 years, allowing for an investigation of the binary fractions in Draco and Ursa Minor. Most recently, 2023A&A...677A..95A applied this approach to measure the binary fraction in the Sculptor dwarf galaxy.
@@
These cumulative efforts highlight the growing interest in dwarf galaxies as valuable targets for binary population studies. The inferred binary fractions vary significantly among different dwarf galaxies, indicating that binary populations are potentially sensitive to the specific conditions of their host systems. Dwarf galaxies themselves exhibit a wide range of structural, chemical, and dynamical properties (e.g. 2012AJ....144....4M; 2024arXiv241107424P). Though dwarf galaxies are mostly old, the star formation histories of different dwarf galaxies still vary (2014ApJ...789..147W). Variations in stellar density, metallicity, and star formation histories among different dwarf galaxies provide a natural laboratory for investigating how these factors influence the formation, evolution, and long-term survival of binary stars.
@@
The Ursa Minor (UMi) dwarf spheroidal galaxy is one of the most well-known satellites of the Milky Way, located at a distance of approximately 76 kpc (1995AJ....110.2131A; 2002AJ....123.3199C). UMi is characterized by an old stellar population with a mean age exceeding 10 Gyr, making it an important target for understanding early galactic evolution and chemical enrichment (2001ApJ...549L..63M; 2010ApJS..191..352K). In recent years, numerous spectroscopic observations of UMi have been conducted, yielding a wealth of LOSV measurements for a large sample of stars in UMi (2003ApJ...588L..21K; 2004ApJ...611L..21W; 2010ApJS..191..352K), which offer a valuable opportunity to probe the binary populations in UMi via the velocity variations.
@@
In this study, we take advantage of the multi-epoch LOSV measurements from the Dark Energy Spectroscopic Instrument (DESI) to estimate the binary fraction in UMi. DESI is one of the most advanced multi-object spectrographs designed for wide-field surveys (2013arXiv1308.0847L; 2016arXiv161100036D; 2016arXiv161100037D; 2022AJ....164..207D; 2023AJ....166..259S). The DESI survey aims to observe more than 40 million galaxies, including targets from the Bright Galaxy Survey (BGS; 2023AJ....165..253H), Luminous Red Galaxies (LRG; 2023AJ....165...58Z), Emission Line Galaxies (ELG; 2023AJ....165..126R) and quasars (2023ApJ...944..107C) to explore the nature of dark energy (2023AJ....165...50M; 2023ApJ...943...68L; 2023AJ....165..124A; 2024AJ....167...62D). Significant cosmological results from DESI’s Early Data Release (EDR; 2024AJ....168...58D), Data Release 1 (DR1; 2025arXiv250314745D) and the future Data Release 2 (DR2) include two-point clustering statistics (2024arXiv241112020D), baryon acoustic oscillations (BAO) signals (2024arXiv240403000D; 2025JCAP...01..124A; 2025PhRvD.112h3514A; 2025PhRvD.112h3515A), galaxy clustering analyses (2024arXiv241112021D), and updated cosmological constraints (2025JCAP...02..021A; 2024arXiv241112022D).
@@
Alongside the cosmological surveys of DESI conducted at dark time with good observing conditions, the DESI Milky Way Survey (MWS) (2023ApJ...947...37C) takes advantage of observing conditions suboptimal for galaxy redshift measurements, such as poor seeing and bright time with significant moonlight contamination. The primary DESI MWS program focuses on stars at Galactic latitudes $|b|>20^{\circ}$, reaching up to a magnitude limit of $r=19$. More than seven million stars are selected from the DESI Legacy Imaging Survey (2019AJ....157..168D) for observation under this program. Its footprint encompasses numerous streams, satellite galaxies, and globular clusters in the Milky Way. So far, the stellar value-added catalogs have been published based on the DESI EDR (2024MNRAS.533.1012K) and DESI DR1 (2025arXiv250514787K), containing over 4 million unique stars with LOSV and stellar parameter measurements. Additionally, the DESI MWS includes secondary targets, such as faint blue horizontal branch (BHB) stars, complementing those observed in the main survey. Beyond these main and secondary targets in the main survey, the DESI MWS also employs “tertiary” dedicated tile programs (2023AJ....165...50M) specifically to observe the M31 galaxy, as well as stellar streams and dwarf satellite galaxies within our Galaxy, including UMi. The unprecedented depth and high-precision measurements provided by DESI allow us to apply this methodology to a significantly larger sample of member stars in UMi. This also enables us to explore potential differences in binary fractions across various stellar populations within UMi.
@@
This paper is structured as follows: In Section 2, we describe the properties of the DESI data used in this work. Section 3 introduces our methodology for constructing the binary simulation and the likelihood function to estimate the binary fraction of a dwarf galaxy. In Section 4, we present the results for the dwarf galaxy UMi and the offset between the metal-rich and metal-poor populations. We discuss the potential causes for this binary fraction discrepancy in Section 5. Finally, the summary and conclusions are provided in Section 6.
2 Data
@@
2.1 DESI
DESI is designed to conduct sky surveys covering 14,000 deg2 on the Mayall 4-meter telescope at the Kitt Peak National Observatory, Arizona (2013arXiv1308.0847L; 2016arXiv161100036D; 2016arXiv161100037D; 2022AJ....164..207D; 2023AJ....166..259S). It is a multi-object spectroscopic instrument capable of collecting up to 5,000 spectra per exposure, providing a continuous wavelength coverage from 3,600 to 9,824 Å  with a full width at half maximum resolution of about 1.8 Å  (2023AJ....165....9S; 2024AJ....168...95M; 2024AJ....168..245P).
In this work, we utilize data collected through DESI MWS tertiary dedicated tile programs 28 and 33. Tertiary 28 observations were conducted on 2023 April 27 while only four of the ten DESI spectrograph petals were in operation. These observations include 12 tiles (i.e., unique fiber configurations) designed such that the inner $\simeq$5 half-light radii of UMi were covered by three adjacent working petals. Tertiary 33 observations were conducted on 7 separate nights between 2023 June 06 and 2024 March 27 while DESI was fully functional. These observations include 24 tiles distributed to 4 pointings such that each tile observes the inner half-light radius of UMi and the full program footprint extended out to $\simeq$10 half-light radii. The footprints of both programs are presented in Figure 1. In both programs, each tile was observed between 350 and 1500 s depending on the observing conditions in order to achieve an effective exposure time of $\simeq$300 s (see 2023AJ....166..259S for a description of how the effective exposure time is calculated). A more detailed description of these UMi tertiary observations will be presented in Sanford et al. (In Prep).
@@ 2.1 DESI
These observations were processed with the latest version of the DESI spectroscopic pipeline(2023AJ....165..144G), yielding 49,836 spectra for 8,460 unique objects. These stellar spectra provide valuable measurements, including LOSVs ($v_{\rm{los}}$), metallicity ([Fe/H]), surface gravity ($\log g$), effective temperature ($T_{\mathrm{eff}}$), a few other chemical abundances, etc., which are output by the MWS rvs [^1] https://github.com/segasai/rvspecfit(2019ascl.soft07013K) and sp pipelines. Readers can refer to 2023ApJ...947...37C for more details about the MWS pipelines. This data will be included as a part of DESI DR2 scheduled to be released in spring 2027.
@@
It is important to note that the uncertainties of LOSV ($\sigma_{v_{\rm{los}}}$) derived from the pipeline have a systematic floor of around 1-2 km s-1, which refers to an irreducible component of measurement uncertainty that persists even when statistical errors are minimized. It typically arises from instrumental effects such as spectrograph flexure or wavelength calibration errors. This floor is modeled as a constant added in quadrature to the formal uncertainty and represents a lower limit to the achievable precision. As discussed in Section 4.3.2 of 2025arXiv250514787K, the systematic error floor varies with the observing condition. The systematic error floors are 1.0, 1.6, and 2.0 km s-1 for the MWS bright, dark, and backup programs (see 2025arXiv250517230D, for details about the MWS backup program), respectively. Given the specific strategy of the tertiary program, we will additionally constrain the systematic error floor for our case to improve the precision of our results in Section 3.5.
2.2 Member Stars in UMi
@@
To obtain a clean sample of member stars in UMi, we match the stars observed by DESI in the UMi footprint with the highly probable (probability greater than 90%) member star lists for MW dwarf spheroidal galaxies selected by 2022ApJ...940..136P [^2] The catalog is available at 10.5281/zenodo.6533295 (andrew_b_pace_2025_14732612). The authors combine astrometry from Gaia early data release 3 (EDR3) and the photometry from either DECam, Gaia, or Pan-Starrs to effectively separate the MW foreground stars from member stars in these dwarf galaxies (2019ApJ...875...77P; 2020AJ....160..124M). They model the proper motions and spatial positions of stars around a satellite with Gaussian mixture models, with different Gaussian components standing for true member stars and MW foreground/background components. In particular, they assume a projected Plummer stellar distribution (1911MNRAS..71..460P) for the spatial term. After matching our data to their catalog, we obtain a sample of 1,721 candidates within the UMi footprint, comprising a total of 8,439 observations. Each member star has multiple observations by DESI at different epochs over the course of a year. The distributions of the Modified Julian Date (MJD) range and the number of epochs per star are shown in Figure 2. For each member star, we adopt the metallicity and surface gravity values derived from the stacked spectra across all exposures and by the MWS rvs pipeline. These values serve as representative parameters for each individual object and will be used in the subsequent selection criteria. Besides, for multiple observations obtained within a one-day interval, we treat them as one single measurement. Specifically, we co-add the original spectra taken on the same night and derive the LOSV and its corresponding uncertainty from the stacked spectrum rather than from individual exposures.
@@
Figure 2: The basic information of our selected member stars (670 in total) of UMi. Top: The histogram showing the MJD distribution for each epoch. Due to the observation strategy of the Tertiary Program, the MJD range for each star spans from several days to almost a year. Bottom: Cumulative histograms of the number of epochs per star. The lighter dashed histogram shows the distribution before stacking multiple spectra taken on the same night (see Step 4 in Section 2.2), while the solid black histogram shows the distribution after merging those exposures into a single epoch.
@@
Figure 3: Top-left panel: Color–magnitude diagram for UMi member stars. The red dashed line marks the color–magnitude selection used to exclude horizontal branch stars. Top-right panel: LOSV distribution for all measurements of the member stars, derived from stacked spectra obtained each night. Bottom-left panel: Spatial distribution of UMi member stars in equatorial coordinates in a gnomonic projection. The red ellipse marks the division between the inner and outer subsamples based on projected radius from the galactic center (see Section 5). The three black ellipses correspond to the 1, 5, and 10 elliptical half-light radii, identical to those shown in Figure 1. Bottom-right panel: Metallicity distribution of UMi member stars, measured from stacked spectra combining all available exposures for each star.
@@
    Stars with exposures all within one day are excluded from the analysis. Our binary fraction estimation relies on detecting variations in LOSV, which requires sufficient time baselines. In this study, we also model the Roche lobe (see Section 3), which effectively excludes binary systems with orbital periods shorter than 4 days. As a result, significant LOSV variations due to binary orbital motions are unlikely to occur within time spans shorter than a quarter of the orbital period, further justifying this exclusion criterion.
@@
In addition, we also match our data with the variable catalog from Gaia DR3 to remove contamination by RR Lyrae. However, there might be some faint RR Lyrae left, which are not identified by Gaia. Thus, we further exclude horizontal branch stars using an empirical cut of $m_{G}>23-5\times(m_{\rm BP}-m_{\rm RP})$. After careful selection above, we have 670 unique member stars spanning over one day, comprising a total of 2,147 individual measurements of LOSV. We summarize these numbers in Table 1. The distributions of basic information about our sample are presented in Figures 2 and 3. As can be seen from the lower left panel, most of our member stars are within 1-1.5 degrees (semi-major axis) to the center, i.e., within 5 times the elliptical half-light radius of UMi (see the middle black ellipse in Figure 1).

### v2

@@ Abstract
thanks: qtskyxt@sjtu.edu.cnthanks: wenting.wang@sjtu.edu.cnthanks: Sergey.Koposov@ed.ac.uk
I Introduction
@@
Binary stars play a crucial role in our understanding of the star formation process, galactic dynamics, and the evolution history of galaxies (White and Ghez, 2001; Duchêne and Kraus, 2013; Badenes et al., 2018). Binaries are common in our Milky Way Galaxy, with studies showing that binary systems are identified more often than single stars in the solar neighborhood (Heintz, 1969; Duquennoy and Mayor, 1991; Raghavan et al., 2010). These earlier studies mostly focus on the statistical properties of binary stars within a limited range, typically extending only to about 30 pc from the Sun.
@@
With advances in observational techniques, investigations of binary populations have extended (Price-Whelan et al., 2018) beyond the solar neighborhood to open clusters (Patience et al., 2002; Li et al., 2020), globular clusters (Pryor et al., 1988; Hut et al., 1992; Sollima et al., 2007; Milone et al., 2012) and dwarf galaxies (Olszewski et al., 1996; McConnachie and Côté, 2010; Minor, 2013). In dense stellar environments, binaries are likely to have undergone significant dynamical processes such as stellar encounters (Kaczmarek et al., 2011; Deacon and Kraus, 2020), which can disrupt binary systems. Compared with globular clusters, dwarf galaxies are more diffuse, making close stellar encounters among their member stars less frequent. The lower stellar densities in dwarf galaxies may allow a higher fraction of binaries to survive than those in dense globular clusters.
@@
Most of the classical dwarf galaxies around our Milky Way are distant, and thus binaries in dwarf galaxies are difficult to resolve directly. The influence of binaries is often inferred from their contribution to observed line-of-sight velocity (LOSV) variations. For example, binary orbital motions can inflate the measured velocity dispersion of a galaxy, thereby complicating dynamical mass estimates (Olszewski et al., 1996; Hargreaves et al., 1996; McConnachie and Côté, 2010; Koposov et al., 2011; Minor et al., 2019; Wang et al., 2023). With advances in spectroscopic surveys, the number of stars observed in dwarf galaxies has increased substantially, from less than a hundred in earlier studies (Queloz et al., 1995; Armandroff et al., 1995; Vogt et al., 1995; Kleyna et al., 2002, 2003) to several hundred or even over a thousand in those known dwarf galaxies such as Carina, Fornax, Sculptor, Sextans, etc (Walker et al., 2009, 2023).
The methodological foundation was established by these early pioneering efforts to search for individual binary systems. Pryor et al. (1988) employed multi-epoch radial-velocity data of giants to identify binary candidates, while Hut et al. (1992) provided a comprehensive review of different binary-detection techniques and their applications in dense stellar systems. Later works extended their efforts beyond individual binary identification to the development of statistical approaches for constraining the overall binary fraction.
@@
In particular, Olszewski et al. (1996) analyzed extensive radial velocities to search binaries and infer their frequency, and demonstrated that the binaries can significantly affect the measured velocity dispersion in dense systems. Based on this points, Minor et al. (2010) introduced a method to estimate the binary fraction in dwarf galaxies based on their velocity dispersion. Simon et al. (2011) applied this approach to Segue 1 to determine the velocity dispersion, removing the influence from the binaries. A slightly later study by Minor (2013) conducted a detailed binary analysis of several dwarf galaxies, including Carina, Fornax, etc. Combining several previous datasets, Spencer et al. (2017b) compiled multi-epoch LOSV measurements and applied a newly developed method to study the binary population in Leo II. A more recent study by Spencer et al. (2018) (hereafter S18) further extended this methodology by combining LOSV measurements from multiple sources collected over approximately 30 years, allowing for an investigation of the binary fractions in Draco and Ursa Minor. Most recently, Arroyo-Polonio et al. (2023) applied this approach to measure the binary fraction in the Sculptor dwarf galaxy.
@@
These cumulative efforts highlight the growing interest in dwarf galaxies as valuable targets for binary population studies. The inferred binary fractions vary significantly among different dwarf galaxies, indicating that binary populations are potentially sensitive to the specific conditions of their host systems. Dwarf galaxies themselves exhibit a wide range of structural, chemical, and dynamical properties (e.g. McConnachie, 2012; Pace, 2024). Though dwarf galaxies are mostly old, the star formation histories of different dwarf galaxies still vary (Weisz et al., 2014). Variations in stellar density, metallicity, and star formation histories among different dwarf galaxies provide a natural laboratory for investigating how these factors influence the formation, evolution, and long-term survival of binary stars.
@@
The Ursa Minor (UMi) dwarf spheroidal galaxy is one of the most well-known satellites of the Milky Way, located at a distance of approximately 76 kpc (Armandroff et al., 1995; Carrera et al., 2002). UMi is characterized by an old stellar population with a mean age exceeding 10 Gyr, making it an important target for understanding early galactic evolution and chemical enrichment (Martínez-Delgado et al., 2001; Kirby et al., 2010). In recent years, numerous spectroscopic observations of UMi have been conducted, yielding a wealth of LOSV measurements for a large sample of stars in UMi (Kleyna et al., 2003; Wilkinson et al., 2004; Kirby et al., 2010), which offer a valuable opportunity to probe the binary populations in UMi via the velocity variations.
@@
In this study, we take advantage of the multi-epoch LOSV measurements from the Dark Energy Spectroscopic Instrument (DESI) to estimate the binary fraction in UMi. DESI is one of the most advanced multi-object spectrographs designed for wide-field surveys (Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022; Schlafly et al., 2023). The DESI survey aims to observe more than 40 million galaxies, including targets from the Bright Galaxy Survey (BGS; Hahn et al., 2023), Luminous Red Galaxies (LRG; Zhou et al., 2023), Emission Line Galaxies (ELG; Raichoor et al., 2023) and quasars (Chaussidon et al., 2023) to explore the nature of dark energy (Myers et al., 2023; Lan et al., 2023; Alexander et al., 2023; DESI Collaboration et al., 2024a). Significant cosmological results from DESI’s Early Data Release (EDR; DESI Collaboration et al., 2024b), Data Release 1 (DR1; DESI Collaboration et al., 2025) and the future Data Release 2 (DR2) include two-point clustering statistics (DESI Collaboration et al., 2024d), baryon acoustic oscillations (BAO) signals (DESI Collaboration et al., 2024f; Adame et al., 2025b; Abdul Karim et al., 2025b, a), galaxy clustering analyses (DESI Collaboration et al., 2024e), and updated cosmological constraints (Adame et al., 2025a; DESI Collaboration et al., 2024c).
@@
Alongside the cosmological surveys of DESI conducted at dark time with good observing conditions, the DESI Milky Way Survey (MWS) (Cooper et al., 2023) takes advantage of observing conditions suboptimal for galaxy redshift measurements, such as poor seeing and bright time with significant moonlight contamination. The primary DESI MWS program focuses on stars at Galactic latitudes $|b|>20^{\circ}$, reaching up to a magnitude limit of $r=19$. More than seven million stars are selected from the DESI Legacy Imaging Survey (Dey et al., 2019) for observation under this program. Its footprint encompasses numerous streams, satellite galaxies, and globular clusters in the Milky Way. So far, the stellar value-added catalogs have been published based on the DESI EDR (Koposov et al., 2024) and DESI DR1 (Koposov et al., 2025), containing over 4 million unique stars with LOSV and stellar parameter measurements. Additionally, the DESI MWS includes secondary targets, such as faint blue horizontal branch (BHB) stars, complementing those observed in the main survey. Beyond these main and secondary targets in the main survey, the DESI MWS also employs “tertiary” dedicated tile programs (Myers et al., 2023) specifically to observe the M31 galaxy, as well as stellar streams and dwarf satellite galaxies within our Galaxy, including UMi. The unprecedented depth and high-precision measurements provided by DESI allow us to apply this methodology to a significantly larger sample of member stars in UMi. This also enables us to explore potential differences in binary fractions across various stellar populations within UMi.
@@
This paper is structured as follows: In Section II, we describe the properties of the DESI data used in this work. Section III introduces our methodology for constructing the binary simulation and the likelihood function to estimate the binary fraction of a dwarf galaxy. In Section IV, we present the results for the dwarf galaxy UMi and the offset between the metal-rich and metal-poor populations. We discuss the potential causes for this binary fraction discrepancy in Section V. Finally, the summary and conclusions are provided in Section VI.
II Data
@@
II.1 DESI
DESI is designed to conduct sky surveys covering 14,000 deg2 on the Mayall 4-meter telescope at the Kitt Peak National Observatory, Arizona (Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022; Schlafly et al., 2023). It is a multi-object spectroscopic instrument capable of collecting up to 5,000 spectra per exposure, providing a continuous wavelength coverage from 3,600 to 9,824 Å  with a full width at half maximum resolution of about 1.8 Å  (Silber et al., 2023; Miller et al., 2024; Poppett et al., 2024).
In this work, we utilize data collected through DESI MWS tertiary dedicated tile programs 28 and 33. Tertiary 28 observations were conducted on 2023 April 27 while only four of the ten DESI spectrograph petals were in operation. These observations include 12 tiles (i.e., unique fiber configurations) designed such that the inner $\simeq$5 half-light radii of UMi were covered by three adjacent working petals. Tertiary 33 observations were conducted on 7 separate nights between 2023 June 06 and 2024 March 27 while DESI was fully functional. These observations include 24 tiles distributed to 4 pointings such that each tile observes the inner half-light radius of UMi and the full program footprint extended out to $\simeq$10 half-light radii. The footprints of both programs are presented in Figure 1. The adopted parameters of UMi used in this work are listed in Table 1. In both programs, each tile was observed between 350 and 1500 s depending on the observing conditions in order to achieve an effective exposure time of $\simeq$300 s (see Schlafly et al. 2023 for a description of how the effective exposure time is calculated). A more detailed description of these UMi tertiary observations will be presented in Sanford et al. (In Prep).
@@ II.1 DESI
Table 1: Summary of adopted parameters of UMi.
Distance 76 kpc
(Carrera et al., 2002)
Half-light radii 18.3 arcmin
(Pace, 2025)
Ellipticity 0.55
(Pace, 2025)
Position angle 50 deg
(Pace, 2025)
Velocity dispersion 8.0 km s-1
Average age for member stars 10 Gyr
(Carrera et al., 2002)
Central luminosity density 0.006 $L_{\odot}\mathrm{pc}^{-3}$
(Mateo, 1998)
@@
These observations were processed with the latest version of the DESI spectroscopic pipeline(Guy et al., 2023), yielding 49,836 spectra for 8,460 unique objects. These stellar spectra provide valuable measurements, including LOSVs ($v_{\rm{los}}$), metallicity ([Fe/H]), surface gravity ($\log g$), effective temperature ($T_{\mathrm{eff}}$), a few other chemical abundances, etc., which are output by the MWS rvs [^1] https://github.com/segasai/rvspecfit(Koposov, 2019) and sp pipelines. Readers can refer to Cooper et al. (2023) for more details about the MWS pipelines. This data will be included as a part of DESI DR2 scheduled to be released in spring 2027.
@@
It is important to note that the uncertainties of LOSV ($\sigma_{v_{\rm{los}}}$) derived from the pipeline have a systematic floor of around 1-2 km s-1, which refers to an irreducible component of measurement uncertainty that persists even when statistical errors are minimized. It typically arises from instrumental effects such as spectrograph flexure or wavelength calibration errors and observing conditions such as lunation fraction, atmosphere, etc. This floor is modeled as a constant added in quadrature to the formal uncertainty and represents a lower limit to the achievable precision. As discussed in Section 4.3.2 of Koposov et al. (2025), the systematic error floors are 1.0, 1.6, and 2.0 km s-1 for the MWS bright, dark, and backup programs (see Dey et al., 2025, for details about the MWS backup program), respectively. Given the specific strategy of the tertiary program, we will additionally constrain the systematic error floor for our case to improve the precision of our results in Section III.5.
II.2 Member Stars in UMi
@@
To obtain a clean sample of member stars in UMi, we match the stars observed by DESI in the UMi footprint with the highly probable (probability greater than 90%) member star lists for dwarf spheroidal galaxies selected by Pace et al. (2022) [^2] The catalog is available at 10.5281/zenodo.6533295 (Pace et al., 2025). The authors combine astrometry from Gaia early data release 3 (EDR3) and the photometry from either DECam, Gaia, or Pan-Starrs to effectively separate the Milky Way foreground stars from member stars in these dwarf galaxies (Pace and Li, 2019; McConnachie and Venn, 2020). They model the proper motions and spatial positions of stars around a satellite with Gaussian mixture models, with different Gaussian components standing for true member stars and foreground/background components. In particular, they assume a projected Plummer stellar distribution (Plummer, 1911) for the spatial term. After matching our data to their catalog, we obtain a sample of 1,721 candidates within the UMi footprint, comprising a total of 8,439 observations. Each member star has multiple observations by DESI at different epochs over the course of a year. The distributions of the Modified Julian Date (MJD) range and the number of epochs per star are shown in Figure 2. For each member star, we adopt the metallicity and surface gravity values derived from the stacked spectra across all exposures and by the MWS rvs pipeline. These values serve as representative parameters for each individual object and will be used in the subsequent selection criteria. Besides, for multiple observations obtained within a one-day interval, we treat them as one single measurement. Specifically, we co-add the original spectra taken on the same night and derive the LOSV and its corresponding uncertainty from the stacked spectrum rather than from individual exposures.
@@
Figure 2: The basic information of our selected member stars (670 in total) of UMi. Top: The histogram showing the MJD distribution for each epoch. Due to the observation strategy of the Tertiary Program, the MJD range for each star spans from several days to almost a year. Bottom: Cumulative histograms of the number of epochs per star. The lighter dashed histogram shows the distribution before stacking multiple spectra taken on the same night (see Step 4 in Section II.2), while the solid black histogram shows the distribution after merging those exposures into a single epoch.
@@
Figure 3: Top-left panel: Spatial distribution of UMi member stars in equatorial coordinates in a gnomonic projection. The red ellipse marks the division between the inner and outer subsamples based on projected radius from the galactic center (see Section V). The three black ellipses correspond to the 1, 5, and 10 elliptical half-light radii, identical to those shown in Figure 1. Top-right panel: Color–magnitude diagram for UMi member stars. The red dashed line marks the color–magnitude selection used to exclude horizontal branch stars. Bottom-left panel: Metallicity distribution of UMi member stars, measured from stacked spectra combining all available exposures for each star. Bottom-middle panel: LOSV distribution for all measurements of the member stars, derived from stacked spectra obtained each night. Bottom-right panel: Magnitude dependence of LOSV uncertainties. Gray points denote individual stars, while black points with errorbars indicate the mean LOSV uncertainty and the corresponding standard error derived in magnitude bins of width 0.25.
@@
    Stars with exposures all within one day are excluded from the analysis. Our binary fraction estimation relies on detecting variations in LOSV, which requires sufficient time baselines. In this study, we also model the Roche lobe (see Section III), which effectively excludes binary systems with orbital periods shorter than 4 days. As a result, significant LOSV variations due to binary orbital motions are unlikely to occur within time spans shorter than a quarter of the orbital period, further justifying this exclusion criterion.
@@
In addition, we also match our data with the variable catalog from Gaia DR3 to remove contamination by RR Lyrae. However, there might be some faint RR Lyrae left, which are not identified by Gaia. Thus, we further exclude horizontal branch stars using an empirical cut of $m_{G}>23-5\times(m_{\rm BP}-m_{\rm RP})$. After careful selection above, we have 670 unique member stars spanning over one day, comprising a total of 2,147 individual measurements of LOSV. We summarize these numbers in Table 2. The distributions of basic information about our sample are presented in Figures 2 and 3. As can be seen from the upper left panel, most of our member stars are within 1-1.5 degrees (semi-major axis) to the center, i.e., within 5 times the elliptical half-light radius of UMi (see the middle black ellipse in Figure 1).